# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

### v1

@@ VI-D Visualization
A visualization of the attention process for the different prediction approaches is provided in Figure VI-D. Displayed images are cropped to a single line for visibility, but predictions are performed on the full-page images. Attention weights of the last cross-attention layer are projected on the input image and values are represented by color intensity (maximum intensity for 0.2 value). One color is associated to each projection head in this order: red, blue, green, yellow and cyan. As one can notice, the attention weights are more peaky for DAN and W-DAN. It was expected as each query only focuses on a single character prediction, contrary to MT-DAN and Meta-DAN that involve a multi-token prediction strategy. Interestingly, the attention weights of the MT-DAN even spread over the last predicted characters ("he " in the example) without disturbing the recognition. The combination of multi-token predictions and window queries enables the Meta-DAN to efficiently predict more tokens (9 here) by taking advantage of more context.
@@
VI-E Large-scale comparison of the prediction approaches
@@
CER results (in percentages) on the test sets with FCN-M encoder. We used $m=5$ and $w=5$ for the proposed approaches. Best results at page level are shown in bold. Dataset Line-level Page-level FCN DAN Faster DAN W-DAN (ours) MT-DAN (ours) Meta-DAN (ours) CER Time CER Time CER Time CER Time CER Time CER Time IAM 4.84 0.03 4.23 3.39 4.16 0.64 4.87 0.77 4.51 0.83 4.79 0.43 BRESSAY 3.45 0.02 2.43 24.65 2.34 1.51 3.02 6.63 2.33 5.70 2.38 4.01 READ-2016 4.61 0.03 3.85 4.88 4.34 0.73 4.34 1.09 3.80 0.80 4.22 0.63 Esposalles 0.61 0.05 5.86 19.99 1.91 2.23 5.95 4.91 2.16 3.37 2.13 2.86 ScribbleLens 4.72 0.06 4.38 16.24 4.81 2.17 5.06 4.04 4.16 2.71 4.31 2.24 CASIA-2 2.91 0.05 1.10 2.81 3.17 0.57 2.81 0.69 1.32 0.52 2.01 0.43 Eparchos 7.79 0.04 7.27 9.14 7.52 0.97 8.18 2.11 8.14 1.52 7.36 1.21 RIMES-2009 6.73 5.87 6.56 0.98 6.26 1.44 4.94 1.09 4.64 0.83 RIMES-2024 3.89 5.89 3.57 0.94 3.48 1.38 2.97 1.06 3.26 0.80 MAURDOR-C3 10.38 5.34 9.76 0.78 8.05 1.27 13.71 1.16 7.93 0.69 MAURDOR-C4 8.28 6.70 10.46 1.32 6.93 1.61 6.62 1.20 6.59 0.92 Average 5.31 9.54 5.33 1.17 5.36 2.36 4.97 1.81 4.51 1.37 # parameters 17.7 M 23.1 M 23.1 M 23.1 M 23.4 M 23.4 M
@@
We now provide a large evaluation at page level over all datasets and all proposed prediction approaches, as well as results for the concurrent Faster DAN decoding strategy. We used the same architecture and training strategy for all the evaluating settings, i.e., we also train the concurrent approaches with the same framework for fair comparisons. We also provide results at line level, which corresponds to the pre-training stage (same encoder with CTC decoding) for comparison. However, let recall that line-level metrics only correspond to recognition on humanly-segmented images: it does not include the required segmentation and ordering stages that would lead to higher CER and prediction times. Results shown in Table VI-E highlight that it is generally beneficial to process documents at page level rather than line level. A counter-example is the Esposalles dataset, where line recognition performs very well, demonstrating a certain ease of visual reading (single writer). In this particular case, it seems that the benefits of implicit language modeling provided by the transformer decoder are not sufficient to compensate for the added difficulty of learning the reading order.
@@
Another point is about the prediction time: the three proposed approaches are on par with the prediction speed of the Faster DAN. However, one can note that the Faster DAN speed improvement is highly dependent to the number of lines to be predicted in the document, contrary to the proposed approaches. Indeed, the Faster DAN is 16.3 times faster than the DAN for the BRESSAY dataset but 4.9 times faster for the CASIA-2 dataset, given that there are on average 30 lines per document for BRESSAY and 10 lines per document for CASIA-2.
@@
Regarding the CER performance now, the MT-DAN approach reaches new state-of-the art performance over 5 datasets. Also, it seems that the enlarged context provided by the Meta-DAN enables to stabilize the attention process, reducing the phenomenon of predictions loops (when the same text is predicted multiple times) for heterogeneous layouts as in the MAURDOC-C3 and MAURDOR-C4 datasets. It results in new state-of-the-art recognition performance when averaging on all datasets with an improvement of 0.46 point compared to MT-DAN and 0.80 point compared to DAN. WER results are given in Appendix C for comparison purposes.
@@
Comparison with state-of-the-art approaches (no external data, no external language model). Best result at page level is shown in bold. State-of-the-art line-level results are given for comparison. [b] Dataset Approaches of the literature CER (%) for our approaches (page level) Architecture Level CER (%) MT-DAN Meta-DAN BRESSAY [42] CRNN + attention line 3.62 2.33 2.38 [42] YOLO-v5 + Calamari line 6.01 [42] DANIEL page 3.55 CASIA-2 [48] CNN + attention line 2.34 1.32 2.01 [49] DenseNet + attention line 1.78 Eparchos [44] CNN + GRU line 4.53 8.14 7.36 Esposalles [21] DAN + NER line 0.48 2.16 2.13 [21] DAN + NER page 3.61 IAM [50] transformer line 3.99 4.51 4.79 [13] DAN page* 4.54 [17] DANIEL page* 4.38 MAURDOR-C3 [13] DAN page 8.62 13.71 7.93 [14] Faster DAN page 8.93 MAURDOR-C4 [13] DAN page 8.02 6.62 6.59 [14] Faster DAN page 9.88 READ-2016 [12] VAN line 4.10 3.80 4.22 [13] DAN paragraph 3.22 [13] DAN page 3.43 [14] Faster DAN page 3.95 [17] DANIEL page 4.03 RIMES-2009 [13] DAN page 4.54 4.94 4.64 [14] Faster DAN page 6.38 [17] DANIEL page 5.80 RIMES-2024 2.97 3.26 ScribbleLens [51] CNN + BLSTM line 6.47 4.16 4.31 * For IAM, the page level refers in the literature to pages for which the header including duplicated printed text was removed.
VI-F Comparison with the state of the art
@@
We further compare the proposed approaches to the state of the art in Table VI-E. For fair comparison, we only compare to works following the same constraints, i.e., without external language model and that do not perform training on additional external dataset. Among the 7 datasets for which there is a page-level evaluation in the literature, we reach new state-of-the-art CER on 4 of these datasets with the MT-DAN or the Meta-DAN. The proposed approaches also outperform state-of-the-art line-level recognition models on ScribbleLens and CASIA-2 datasets. On can note an exception with the Eparchos dataset, which follow a similar behavior compared to the Esposalles dataset, i.e., text recognition is better at line level than at page level. This can be explained by the very few available training data that prevent from fully benefiting from the full potential of attention-based recognition systems. Lastly, we provide a first evaluation for the newly released RIMES-2024 dataset with a CER of 2.97% reached by the MT-DAN.
VI-G Multilingual recognition
CER (in percentages) for different encoders on the test sets for line-level multilingual recognition. FCN-M ConvNeXtV2-N ConvNeXtV2-B Pre-training IN-21k IN-21k IN-21k Training Monolingual Multilingual Monolingual Multilingual Multilingual Multilingual IAM 4.84 6.40 5.54 9.57 6.34 5.31 BRESSAY 3.45 5.27 3.61 7.85 4.51 3.23 READ-2016 4.61 6.92 5.23 8.28 6.25 4.74 Esposalles 0.61 1.83 1.02 2.11 1.33 0.82 ScribbleLens 4.72 6.49 5.27 8.61 6.07 5.07 Average CER (%) 3.65 5.38 4.13 7.28 4.90 3.83 # parameters 17.6 M 16.0 M 88.8 M Average time (s) 0.03 0.01 0.03
In this last experiment, we propose training the Meta-DAN in a multilingual setting, i.e., the model is trained on a set of datasets at once, and evaluated on the same datasets. We only focus on latin languages to see if learning one language can be helpful for learning another one when sharing part of the same alphabet.
We compare the FCN-M encoder we used so far to two variants of the ConvNeXt-V2 architecture [52], which is a state-of-the-art CNN architecture for computer vision, pre-trained through self-supervised learning (masked auto-encoder) on ImageNet-21k (IN-21k). Regarding these ConvNeXt-V2 encoders, we made the following modifications for fair comparison and CTC requirements for pre-training (minimum input length with respect to the ground truth): 1) the last projection is set to output 256 feature maps to be compatible with the DAN decoder; 2) The ConvNeXt encoder includes five downsampling stages, each one dividing by two the horizontal and vertical axes. We removed two of them for the horizontal axis, leading to a final output shape for the image features of ($H_{f}=\frac{H}{32},W_{f}=\frac{W}{8},C_{f}=256$), which corresponds to the FCN-M encoder output shape.
We first concentrate on the line-level pre-training stage in Table VI-G. As one can note, the FCN-M encoder is better suited for monolingual setting compared to the same-scale ConvNeXt-V2-N model, but we can observe the contrary for the multilingual setting. The results highlight that is mainly due to the self-supervised pre-training strategy on a large-scale IN-21k dataset: the overall CER is 2.38 point higher when training from scratch. Another interesting point is that, even with a larger model (ConvNeXt-V2-B, 4 times more parameters than FCN-M), the obtained results are still a bit less good as the monolingual approach with FCN-M encoder (3.83% of CER compared to 3.65%).
We now compare these different encoders at page level with the Meta-DAN approach. To make training more efficient, we add an extra training step at line-level, leading to a 4-step training: 1) pre-training the encoder with CTC supervision through the line-level approach; 2) transfer learning to MT-DAN on line-level images; 3) further training the MT-DAN at page level with curriculum learning 4) transfer learning to Meta-DAN at page level without curriculum strategy.
Table VI-G describes the obtained results, highlighting the need for an upscaled encoder to get close to monolingual results. Indeed, the ConvNeXt-V2-B encoder leads to an average CER of 4.62% on a multilingual setting, which is very close to the 4.45% obtained by the FCN-M with monolingual training. In addition, it only results in a small prediction time increase, from 1.57 second per page (FCN-M, monolingual) to 1.74 second per page (CN2-B, multilingual).
Comparison of CER results (in percentages) at page-level with FCN-M and ConvNeXt-V2 (CN2) encoders for multilingual training using the Meta-DAN ($w=5$, $m=5$). Dataset Monolingual Multilingual FCN-M FCN-M CN2-N CN2-B BRESSAY 2.38 4.82 4.14 4.07 READ-2016 4.22 5.43 6.80 4.91 Esposalles 2.13 2.58 2.69 2.47 ScribbleLens 4.31 9.79 7.51 5.87 IAM 4.79 5.10 4.89 3.97 RIMES-2024 3.26 3.22 3.31 2.93 MAURDOR-C3 7.93 7.41 6.91 7.84 MAURDOR-C4 6.59 17.94 7.06 4.92 Average CER 4.45 7.04 5.41 4.62 Average time (s) 1.57 1.36 1.01 1.74 # parameters 23.2 M 23.2 M 21.5 M 94.3 M
These results also demonstrates that the multilingual setting is only beneficial for similar languages regarding the evaluated models. Comparing monolingual FCN-M and multilingual CN2-B, one can notice that CER of all French/English datasets (IAM, RIMES-2024 and MAURDOR) are improved, while all datasets of historical documents (Esposalles, READ-2016 and ScribbleLens), which involves a specific period and language, are degraded.
VII Discussion
@@
Through this extensive study on prediction strategies, we showed the robustness of the proposed Meta-DAN approach through many evaluation configurations (various datasets, multilingual recognition), reaching on-average state-of-the-art results. However, one can note that there is no one strategy that is better than the others on all datasets. We noticed that, no matter the prediction variant (Faster DAN or ours), it seems to have a beneficial impact on the loop effects that can occur during decoding, noticeable through the CER results of the Esposalles dataset (Table VI-E). We assume that for the Faster DAN strategy, this is due to the parallelization of the line recognition. Indeed, the loops mainly occur when switching from one line to another. For the approaches we proposed, this must be the result of the increased context modeling and the forced anticipation of the model to predict the following characters, which can be found on two consecutive lines. But contrary to the Faster DAN, the Meta-DAN does not suffer from "catastrophic" errors. Indeed, any missed or duplicated character in the first decoding stage of the Faster DAN lead to discard or duplicate a whole text line. While rare, this can severely impact the recognition performance, also explaining the recognition gap. Although we limit ourselves to compare the approaches for character-level predictions to reach state-of-the-art results with few training data, it has to be noted that all approaches could also be used in combination with BPE or N-gram tokenization to further reduce the prediction time.
VIII Conclusion
@@
In this paper, we proposed and evaluated two new prediction strategies for end-to-end full-page text recognition: windowed queries and multi-token predictions. We show that the Meta Document Attention Network approach, combining both strategies, reaches new on-average state-of-the-art results on many handwritten datasets while reaching competitive prediction times. In addition, this approach enables flexible balance between prediction time and recognition performance by adapting the associated hyperparameters to adapt to the use case. While we only evaluated the Meta-DAN for HTR, the proposed approach is generic enough to be adapted to any autoregressive transformer-based system. Lastly, we highlight that using a generic self-supervised pre-trained encoder is an effective way to bridge the performance gap between monolingual and multilingual settings.
@@ Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011012155R2 made by GENCI. Experiments presented in this paper were also partly carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr).
APPENDIX
VIII-A Scaling up the DAN encoder
We noticed that the DAN lacked of some feature extraction capability when evaluated on challenging datasets such as BRESSAY which presents some crossed-out text with annotations. We opt for scaling up the DAN encoder only in order to increase the performance while preventing a significant prediction time increase. Indeed, the encoding part of the model is only used once at inference (it is excluded from the iterative decoding stage); so the computation involved are not such significant compared to decoding. The proposed encoder is referred to as FCN-M, and the original one as FCN-S.
The proposed modification mainly consists in doubling the number of feature maps per block while preserving the last one which condition the size of the decoder, as detailed in Table VIII-A. We use the notation of the original paper [12] where ConvBlock is a Convolution Block made up of three convolutions, instance normalization and dropout, and DSCBlock is equivalent except for the convolutions that are replaced by Depthwise Separable Convolutions [36].
Composition of the two versions of the encoder. ConvBlock(X) and DSCBlock(X) indicate the number of feature maps X returned by each convolutional layer. FCN-S [12] FCN-M (ours) Output size layers ConvBlock(16) ConvBlock(32) $H\times W$ ConvBlock(32) ConvBlock(64) $\frac{H}{2}\times\frac{W}{2}$ ConvBlock(64) ConvBlock(128) $\frac{H}{4}\times\frac{W}{4}$ ConvBlock(128) ConvBlock(256) $\frac{H}{8}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{16}\times\frac{W}{8}$ ConvBlock(128) ConvBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(128) DSCBlock(512) $\frac{H}{32}\times\frac{W}{8}$ DSCBlock(256) DSCBlock(256) $\frac{H}{32}\times\frac{W}{8}$ params 1.7 M 17.6 M
We evaluate these modifications on READ-2016, CASIA-2, BRESSAY and IAM datasets at line level, and at page level with the DAN. Results are shown in Table VIII-A. As one can note, the modified encoder improves the recognition results (or is equivalent) on all the evaluated datasets, whether it is at line or page level. The prediction time is only slightly impacted, we assume this is mainly due to the complexity of learning the reading order. For CASIA-2 and IAM, the prediction time is lower, which can be explained by a reducing of the loop effect (predicting several times the same text part). But for READ-2016 and BRESSAY, the prediction time is higher, which can be the result of the additional computations due to the bigger encoder, and to a better recognition (missed text parts are now well recognized, adding extra decoding iterations). The most negative impact is on the READ-2016 dataset with a 6% increase of the prediction time. These results also enhance what have been found in [13]: the page-level DAN outperforms its line-level counterpart while they are based on the same encoder.
VIII-B Datasets
Line-level splits
Table VIII-B details the split for training, validation and test used for the different datasets regarding the experiments at line level.

### v2

@@ 6.4 Visualization
A visualization of the attention process for the different approaches is provided in Figure 6.4. Images are cropped to a single line for visibility, but predictions are performed on the full-page images. Attention weights of the last cross-attention layer are projected on the input image (values are normalized such that the maximum intensity is reached for 0.2). One color is associated with each projection head in this order: red, blue, green, yellow, and cyan. As expected, the attention weights are peakier for DAN and W-DAN since each query only focuses on a single character prediction, contrary to MT-DAN and Meta-DAN that involve a multi-token prediction strategy. Interestingly, the attention weights of the MT-DAN even spread over the last predicted characters ("he ") without disturbing the recognition. The combination of multi-token predictions and window queries enables the Meta-DAN to efficiently predict more tokens (9 here) by taking advantage of more context.
@@
6.5 Large-scale comparison of the prediction approaches
@@
CER results (in percentages) on the test sets with FCN-M encoder. We used $m=5$ and $w=5$ for the proposed approaches. Best results at page level are in bold. Dataset Line-level Page-level FCN DAN Faster DAN W-DAN MT-DAN Meta-DAN CER Time CER Time CER Time CER Time CER Time CER Time IAM 4.84 0.03 4.23 3.39 4.16 0.64 4.87 0.77 4.51 0.83 4.79 0.43 BRESSAY 3.45 0.02 2.43 24.65 2.34 1.51 3.02 6.63 2.33 5.70 2.38 4.01 READ-2016 4.61 0.03 3.85 4.88 4.34 0.73 4.34 1.09 3.80 0.80 4.22 0.63 Esposalles 0.61 0.05 5.86 19.99 1.91 2.23 5.95 4.91 2.16 3.37 2.13 2.86 ScribbleLens 4.72 0.06 4.38 16.24 4.81 2.17 5.06 4.04 4.16 2.71 4.31 2.24 CASIA-2 2.91 0.05 1.10 2.81 3.17 0.57 2.81 0.69 1.32 0.52 2.01 0.43 Eparchos 7.79 0.04 7.27 9.14 7.52 0.97 8.18 2.11 8.14 1.52 7.36 1.21 RIMES-2009 6.73 5.87 6.56 0.98 6.26 1.44 4.94 1.09 4.64 0.83 RIMES-2024 3.89 5.89 3.57 0.94 3.48 1.38 2.97 1.06 3.26 0.80 MAURDOR-C3 10.38 5.34 9.76 0.78 8.05 1.27 13.71 1.16 7.93 0.69 MAURDOR-C4 8.28 6.70 10.46 1.32 6.93 1.61 6.62 1.20 6.59 0.92 Average 5.31 9.54 5.33 1.17 5.36 2.36 4.97 1.81 4.51 1.37 # parameters 17.7 M 23.1 M 23.1 M 23.1 M 23.4 M 23.4 M
@@
We now provide a large evaluation at page level over all datasets and prediction approaches (ours and concurrent ones). We used the same architecture and training strategy for all the evaluated settings, i.e., we also train the concurrent approaches with the same framework for fair comparisons. We also provide results at line level, which corresponds to the pre-training stage (same encoder with CTC decoding) for comparison. However, let us recall that line-level metrics only correspond to recognition on humanly-segmented images: it does not include the required segmentation and ordering stages that would lead to higher CER and prediction times. Results shown in Table 6.5 highlight that it is generally beneficial to process documents at page level rather than line level. A counter-example is the Esposalles dataset, where line recognition performs very well, demonstrating a certain ease of visual reading (single writer). In this particular case, it seems that the benefits of implicit language modeling provided by the transformer decoder are not sufficient to compensate for the added difficulty of learning the reading order.
@@
Focusing on the prediction time, all proposed approaches are on par with the prediction speed of the Faster DAN. However, one can note that the Faster DAN speed improvement is highly dependent on the number of lines to be predicted in the document, contrary to the proposed approaches. E.g., the Faster DAN is 16.3 times faster than the DAN for the BRESSAY dataset ($\approx 30$ lines/doc) but 4.9 times faster for the CASIA-2 dataset ($\approx$ 10 lines/doc).
@@
Regarding the CER, the MT-DAN reaches state-of-the-art performance on 4 datasets. Also, it seems that the enlarged context provided by the Meta-DAN stabilizes the attention process, reducing the looping phenomenon (when the same text is predicted multiple times) for heterogeneous layouts as in the MAURDOR datasets. It results in new state-of-the-art recognition performance when averaging on all datasets with an improvement of 0.46 points compared to MT-DAN and 0.80 points compared to DAN. Associated WER results are given in C for completeness.
6.6 Comparison with the state of the art
@@
Comparison with state-of-the-art approaches (no external data, no external language model). Best result at page level is shown in bold. State-of-the-art line-level results are given for comparison. [b] Dataset Approaches of the literature CER (%) for our page-level approaches Architecture Level CER (%) MT-DAN Meta-DAN BRESSAY [17] CRNN + attention line 3.62 2.33 2.38 [17] YOLO-v5 + Calamari line 6.01 [17] DANIEL page 3.55 CASIA-2 [47] CNN + attention line 2.34 1.32 2.01 [44] DenseNet + attention line 1.78 Eparchos [42] CNN + GRU line 4.53 8.14 7.36 Esposalles [40] DAN + NER line 0.48 2.16 2.13 [40] DAN + NER page 3.61 IAM [18] transformer line 3.99 4.51 4.79 [14] DAN page 4.54 [11] DANIEL page 4.38 MAURDOR-C3 [14] DAN page 8.62 13.71 7.93 [16] Faster DAN page 8.93 MAURDOR-C4 [14] DAN page 8.02 6.62 6.59 [16] Faster DAN page 9.88 READ-2016 [15] VAN line 4.10 3.80 4.22 [14] DAN page 3.43 [16] Faster DAN page 3.95 [11] DANIEL page 4.03 RIMES-2009 [14] DAN page 4.54 4.94 4.64 [16] Faster DAN page 6.38 [11] DANIEL page 5.80 RIMES-2024 2.97 3.26 ScribbleLens [20] CNN + BLSTM line 6.47 4.16 4.31
@@
We further compare the proposed approaches to the state of the art in Table 6.6. For fair comparison, we only compare to works following the same constraints, i.e., without an external language model and that do not perform training on additional external datasets. Among the 7 datasets for which there is a page-level evaluation in the literature, we reach new state-of-the-art CER on 4 of these datasets with the MT-DAN or the Meta-DAN. The proposed approaches also outperform state-of-the-art line-level recognition models on ScribbleLens and CASIA-2 datasets. One can note an exception with the Eparchos dataset, which follows a similar behavior compared to the Esposalles dataset, i.e., text recognition is better at line level than at page level. This can be explained by the very few available training data that prevents from fully benefiting from the potential of attention-based models.
6.7 Error analysis
In this section, we propose a qualitative analysis of the errors made by the Meta-DAN. To this end, some examples of predictions and associated ground truths are shown in Figure 6.7 to illustrate the different kinds of errors that can occur.
The most common error is linked to visual ambiguities. For instance, the Meta-DAN predicted an "m" instead of "rr" for the word "arrived" (Figure 6.7) or and "n" instead of an "r" in the word "garida" (Figure 6.7). This shows that, although the proposed contributions improve context modeling, it remains limited: the model sometimes tends to rely more on visual characteristics than on linguistic context. We assume that this is due to the lack of training data in the datasets that prevents the model from really "understand" the text it predicts, leading to the prediction of "dest" instead of "de St" (Figure 6.7), due to the spacing between those two words that is not clearly visible in the image. In other words, it suffers from a lack of linguistic knowledge, but it also lacks knowledge of the world around us to clear up certain ambiguities.
Another kind of error comes from the attention-based prediction process itself. In this way, the same character might be predicted twice (Figure 6.7), or some characters are missed, as in Figure 6.7 ("P.S"), generally due to specific layouts not seen during training. In rare cases, some characters may have been predicted in the wrong order and partially missed, as "Davenun" predicted instead of "Dans un avenir" (Figure 6.7). As mentioned previously, another failure case due to the attention mechanism is related to loops, when the model predicts the same part of text, generally a whole text line, multiple times.
Although rare, predictions may be considered false, wrongly, when the ground truth contains errors as in Figure 6.7: the Meta-DAN predicted "d’en", as visible in the input image, but the ground truth is "de".
CASIA-2
IAM
Esposalles
MAURDOR-C3
RIMES-2009
Visualization of errors made by Meta-DAN on test samples of various datasets. For visibility, page images were cropped and associated with Ground Truth (GT) and Prediction (P). Errors are underlined in red.
7 Discussion
@@
Through this extensive study on prediction strategies, we showed the robustness of the proposed Meta-DAN on many HTR dataset configurations, reaching on-average state-of-the-art results. However, one can note that there is no one strategy that is better on all datasets. We noticed that, no matter the prediction variant, it seems to have a beneficial impact on the loop effects, noticeable through the CER results of the Esposalles dataset (Table 6.5). We assume that for the Faster DAN, this is due to the parallelization of the line recognition, since loops mainly occur when switching from one line to another. For the proposed approaches, this must be the result of the increased context modeling and the forced anticipation of the model to predict the following characters, which can be found on two consecutive lines. But contrary to the Faster DAN, the Meta-DAN does not suffer from "catastrophic" errors. Indeed, any missed or duplicated character in the first decoding stage of the Faster DAN leads to discarding or duplicating a whole text line. While rare, this can severely impact the recognition performance.
Regarding the prediction time, it has to be noted that the Meta-DAN, while competitive, does not outperform the Faster DAN in this respect, on average. Indeed, although the hyperparameters $m$ and $w$ can be arbitrarily chosen, the experiments showed that in practice, they must be set to a rather low value to preserve competitive recognition performance. Thus, the current state-of-the-art prediction speed for end-to-end page-level text recognition remains an issue for real-time applications. Although we limit ourselves to comparing the approaches for character-level predictions to reach state-of-the-art results with few training data, all approaches could be combined with BPE tokenization to further reduce the prediction time.
8 Conclusion
@@
In this paper, we proposed two new prediction strategies for end-to-end full-page text recognition: windowed queries and multi-token predictions. We show that the Meta Document Attention Network approach, combining both strategies, reaches new on-average state-of-the-art results on many handwritten datasets while reaching competitive prediction times. This approach offers a flexible balance between prediction time and recognition performance by adapting the associated hyperparameters to the given use case. While we only evaluated the Meta-DAN for HTR, the proposed approach could be applied to any autoregressive transformer-based system.
However, there is still room for improvement, and some issues remain to be solved regarding end-to-end full-page text recognition. The evaluation protocol of state-of-the-art approaches is based on training a specific model on a single dataset at a time. How to scale them to handle multilingual content, multiple reading orders (mixing Latin and Arabic languages, for instance), and highly heterogeneous layouts remains to be explored. These limited training configurations prevent the models from acquiring large, generic linguistic knowledge, leading to some failure cases when the visual content is degraded or ambiguous. Another point is about the prediction time, which remains an issue for real-time applications, as it is still in the order of magnitude of one second per document. Finally, current state-of-the-art approaches consider a single reading order per document, but this is not realistic for very complex layouts (e.g., schemas, maps), for which several reading orders could be considered as humanly correct. Future works should focus on adapting metrics and losses to allow multiple reading orders for those cases.
@@ Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011012155R2 made by GENCI. Experiments were also carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr).
Appendix A Datasets